OpenStudy (anonymous):

Rewrite each expression without exponential expression: 1.) 16^3/4 2.) x^3/5 3.) (4x+2)^3/2 4.) 1/x^2/3

OpenStudy (anonymous):

1) 2 2) \[\sqrt[5]{x^{3}}\] 3) \[(4x + 2)\sqrt{(4x + 2)}\] 4) \[\sqrt[3]{(1/x)^2}\]